OpenStudy (anonymous):

The reaction A+B-> C+D rate=k[A][B]^2 has an initial rate of 0.0810 M/s. What will the initial rate be if [A] is halved and [B] is tripled? What will the initial rate be if [A] is tripled and [B] is halved?

OpenStudy (anonymous):

lets say when the A is hlaved and B is tripled, [1/2A] = 1/2 [A] and [3B]^2 = 9[B] therefore the new rate = k*(1/2[A]) * ( 9[B]) = ( 9/2)k [A][B]. hence the new rate increase by a factor of 9/2 if A tripled and B halved, rate = k * (3[A]) * (1/4 [B]^2) rate = (3/4)k [A][B]. Hope this helps =)